OpenStudy (mokeira):

someone please explain to me the basics of how to graph solutions of inequalities on a number line.

OpenStudy (campbell_st):

ok.... so the basics are solid circle on the number shows the inequalities \[\ge, .. or.. , \le \]. It means include the starting number then an arrow in the same direction as the solution  the example shows \[x \ge 6\] open circles mean don't include and are used for > or <  they are the basics... compound inequalities you can have 2 arrows... on different directions or an interval betweeen numbers hope it helps